Attendance for today's session: Cade, Julie, Alex M, Alex P, Mitchell, Scott.
Monday was a holiday, and we will be replaceing that meeting later this week. Today we worked on the following tasks:
Tasks:
- Check continuity of loop
- Alex M will try another run tonight, and has already requested a virtual desktop.
- Install Arasim to /fs/projects/PAS0654/BiconeEvolutionOCS to be edited and messed with
- Cade is testing whether or not this is running within the loop. He had gotten a few errors, and it is not completing. It's submitting the job with the new bash script, but it is not actually running.
- Found an error. The bash script was not grabbing the correct AraSim directory. More updates to come.
- Clean bash script
- Mitchell and Alex M are putting all current functions hard-coded in into their own .sh scrips that are now being called in the main bash script.
- Update Manual
Right now Elliot and Leo have decided to focus on working with Keith. They will not be contributing to this project for the remainder of the semester. Students have been posting individual updates here: https://www.dropbox.com/home/GP_Antennas/Updates
I've also attached the meeting minutes from the call last week, in order to keep track of updates. |